# Synergizing Foundation Models and Federated Learning: A Survey

### v1

@@ 4.3.2 Privacy Protection
Privacy attacks in FM-FL involve extracting sensitive information from the data used in training, even though the data itself is not directly shared. Major attacks include membership inference attack and data reconstruction attack, where the former aims to determine whether a specific data sample is in a victim client’s training set, and the latter strives to reconstruct original input data from the model parameters or gradients Ren et al. (2024). Regarding membership inference attacks, Vu et al. (2024) revealed the vulnerabilities of popular LLMs, including BERT, DistilBERT, and OpenAI’s GPTs. In terms of data reconstruction attacks, Gupta et al. (2022) presented an attack FILM, which recovers private text data by extracting information from gradients transmitted during training despite employing a DP mechanism.
4.3.3 Attack Robustness
@@
Due to the distributed characteristic of optimization, FL is vulnerable to poisoning attacks Lyu et al. (2022); Rodríguez-Barroso et al. (2023), wherein certain participants may deviate from the prescribed update protocol and upload arbitrary parameters to the central server.
Poisoning Attacks
Depending on the adversarial goals, poisoning attacks in FL can be classified as targeted and untargeted Jere et al. (2020). Targeted attacks, like backdoor attacks, aim to manipulate the global model to generate attacker-desired misclassifications for some particular samples Xie et al. (2020); Bagdasaryan et al. (2020). In contrast, untargeted attacks seek to degrade the model’s overall performance indiscriminately Fang et al. (2020). In addition to the well-recognized attacks on conventional FL studies Li et al. (2023b, 2024b), FM-FL also faces potential threats from compromised pre-trained FMs Li et al. (2023c). Thus, The attacker can introduce backdoors to downstream tasks without prior knowledge Shen et al. (2021). Specifically, Li et al. (2023d) proposed Fed-EBD that introduces a backdoor-compromised FM to generate a public, synthetic dataset for FL training. The clients’ models, pre-trained on this dataset, inherit the backdoor throughout the training.
Defense Techniques
@@
As for defenses, robust aggregation rules are widely applied to make an attack-resilient estimation of the true updates and exclude the influence of malicious updates Blanchard et al. (2017); Yin et al. (2018); Chen et al. (2017); Li et al. (2023a). Other research directions include trust-based strategies Cao et al. (2021); Xu et al. (2022); Park et al. (2021) and variance-reduced algorithms Gorbunov et al. (2023); Wu et al. (2020b). Although these techniques have been widely examined in various FL settings, their effectiveness has yet to be explored in the FM-FL paradigm.
@@
Table 2: A list of representative studies on the applications of FM-FL. Abbreviations: LoRA Tuning (LT), Adapter Tuning (AT), Full-Parameter Tuning (FT), Selective Tuning (ST), Prompt Tuning (PT).
@@
Domain/Application Task Representative Work On-Device Personalization Modality Backbone Fine-Tuning Multilingual NLP Language Understanding FedKC Wang et al. (2022) Txt. mBERT FT Multi-Tasks PMMFL Weller et al. (2022) Txt. mBERT FT Machine Translation Fed-MNMT Liu et al. (2023d) Txt. mBART-50 AT Machine Translation FL-MetaSend Chu et al. (2024) Txt. M2M-100 ST Multi-Tasks MFPT Zhao et al. (2024b) Txt. XLM-RoBERTa PT Speech Speech-to-Text pFedS2T Du et al. (2024) Aud. Conformer/Whisper LT Speech Recognition FedASR Jia et al. (2023) Aud. RNN-T AT Speech Recognition FedE2EASRAzam et al. (2023a) Aud. CTC-AED FT Recommendation General PPLR Zhao et al. (2024a) Txt. LLaMA-7B/LongFormer FT General TransFR Zhang et al. (2024a) Txt. DistBERT AT General GPT-FedRec Zeng et al. (2024) Txt. ChatGPT NA Healthcare Mental Health Prediction FedTherapist Shin et al. (2023a) Txt. BERT & LLaMa-7B LT MRI Reconstruction FedPR Feng et al. (2023a) Vis. Swin Transformers PT
5 Applications of FM-FL
@@
In this part, we briefly review the recent progress on FM-FL applications. Table 2 lists representative work on specific applications and domains.
5.1 FM-FL for Multilingual NLP
@@
Multilingual NLP refers to the techniques that handle multiple natural languages Pires et al. (2019), often to perform equally well across them Wu and Dredze (2020). Earlier research Johnson et al. (2017) has shown that parameter sharing among different languages boosts the model’s performance in multilingual NLP, especially for low-resource languages for which significantly less content is available. However, real-world multilingual text data is often distributed across devices or regions, with each client (user) accessing only a limited subset of languages, where transferring the data to a central server is often problematic or prohibited due to privacy issues Wang et al. (2022). Thanks to its inherent privacy-preserving characteristic, FL holds promise in breaking the barriers of cross-lingual modeling and data isolation by allowing models to learn from decentralized datasets.
@@
The pioneer work by Weller et al. (2022) has firstly demonstrated that fine-tuning pre-trained language models with FL can perform similarly to pre-trained models fine-tuned with the standard centralized method under multilingual NLP settings. Various subsequent studies have focused on adapting pre-trained FMs through FedPEFT techniques such as adapter tuning Liu et al. (2023d), prompt tuning Zhao et al. (2024b), and LoRA Guo et al. (2024b), aiming to enhance training efficiency.
@@
Considering the adverse effect of conflicting parameters from diverse languages during federated fine-tuning, recent studies have exploited clustering strategies to alleviate this issue. For instance, Wang et al. (2022) applied $k$-means clustering on each client’s data to obtain representative knowledge, specifically the clustered data centroids. These centroids were then shared across clients for local training, enriching training data and addressing the challenges associated with data heterogeneity. Another compelling strategy along this line is language family-based clustering. Liu et al. (2023d) explored various clustering strategies to group adapter parameters to mitigate the negative effects of multilingual data heterogeneity, showing that language family-based clustering significantly outperforms the other clustering strategies. Similarly, Guo et al. (2024b) proposed fine-tuning FMs with LoRA and language family-based clustering to address the heterogeneity issue of multilingual modeling.
@@
General downstream tasks include language modeling Wang et al. (2022), machine translation Liu et al. (2023d); Chu et al. (2024), and text classification Weller et al. (2022). In addition, some studies also focus on more specific applications such as medical transcript analysis Manoel et al. (2023) and hate speech detection Akshay and Rahul (2024). These advancements illustrate the applicability of FM-FL across a wide range of scenarios in multilingual NLP.
5.2 FM-FL for Speech
@@
With the development of AI, researchers have also carried out many studies on speech-related FMs, e.g., wav2vec 2.0 Baevski et al. (2020) and Whisper Radford et al. (2023). In this field, the adaptation of FMs often relies on FL to facilitate scenarios where the audio data is privacy-sensitive. Compared to other data modalities, speech-related FM-FL applications especially attract excessive attention to the aspects of on-device training and personalization, motivated by the following considerations: (1) Audio data is continually generated on end-devices such as mobile phones, and owned by individual users—thus it should be processed locally, rather than being transferred elsewhere; (2) Although FL takes advantage of all user data to collectively train one model that maximizes speaker-independent accuracy, such a one-model-fits-all solution can be sub-optimal for individual users Jia et al. (2023). Specific tasks in this field include Automatic Speech Recognition (ASR) Azam et al. (2023b) and Speech-to-Text (S2T) Du et al. (2024).
5.3 FM-FL for Recommendation
@@
Federated Recommendation (FR) strives to capture underlying user preferences and recommend appropriate information to users while safeguarding data privacy Bobadilla et al. (2013); Zhang et al. (2023a). Typical FR systems consist of a server and multiple clients, where clients represent individual users or local data servers possessing smaller datasets and retaining private user information Ammad-Ud-Din et al. (2019). These clients collaborate to train a global model while ensuring their data privacy protection by abstaining from direct data sharing Zeng et al. (2024); Zhang et al. (2023a). Recently, LLM-based recommendations have been gaining increasing attention Wu et al. (2023b) due to their strong capacities in language understanding and domain generalization. The benefits are mainly twofold: (1) LLMs mitigate the cold-start issue by utilizing textual descriptions to make recommendations without the need for extensive historical data Zhang et al. (2023c); (2) The inherent transferability of LLMs allows them to apply cross-domain knowledge and side information to improve accuracy and relevance across diverse items and user interests Gao et al. (2023).
@@
One straightforward way to adapt FMs for FR is by fine-tuning them with historical user-item data. More specifically, FedPEFT techniques such as adapter tuning Zhang et al. (2024a) and split learning Zhao et al. (2024a) can be employed to improve resource efficiency. Apart from parameter fine-tuning, LLMs can also be adapted to assist the recommendation in a zero-shot paradigm through prompt engineering (i.e., without parameter tuning) Gao et al. (2023). For example, Zeng et al. (2024) proposed GPT-FedRec, a two-stage FR framework that leverages ChatGPT for its powerful zero-shot generalization ability. Firstly, GPT-FedRec facilitates hybrid retrieval by collaboratively training ID and text retrievers, after which the retrieved results are transformed into text prompts and submitted to GPT for re-ranking in the second stage. Additionally, Guo et al. (2024a) employed a pre-trained BERT to obtain the representation vectors of item descriptions, which are then fed into a recommender system as augmented input.
5.4 FM-FL for Healthcare

### v2

@@ 6.2.1. Privacy Attacks
Privacy attacks involve extracting sensitive information from the data used in training, even though the data itself is not directly shared. Major attacks include membership inference attack and data reconstruction attack, where the former aims to determine whether a specific data sample is in a victim client’s training set, and the latter strives to reconstruct original input data from the model parameters or gradients (Ren et al., 2025).
Membership Inference Attack.
Vu et al. (2024) design two active membership inference attacks with guaranteed theoretical success rates to assess the privacy leakages of FM fine-tuning under FL configurations, where the FC-based attack activates specific neurons only when a target pattern is present to reveal membership through gradient updates, and the attention-based attack manipulates self-attention heads to filter out the target token, exposing its presence by comparing attention outputs. Similarly, Wen et al. (2024) introduce a privacy backdoor attack, where an adversary poisons a pre-trained FM to amplify privacy leakage during fine-tuning, making membership inference attacks significantly more effective by modifying model weights to induce anomalous loss patterns on targeted data points.
Data Reconstruction Attack.
Gupta et al. (2022) present an attack FILM, which recovers private text data by extracting information from gradients transmitted during FL despite employing a DP mechanism. The attack involves extracting a bag of words, reconstructing sentences via beam search, and refining them with a reordering strategy. Chu et al. (2023) introduce a “panning” attack, which allows a malicious server to extract specific sequences of text containing sensitive information (e.g., credit card numbers) from aggregated user updates, even under large-scale aggregation. The key idea is the use of maliciously modified model parameters to filter and encode relevant sequences in the gradient updates.
6.2.2. Privacy-Preserving Techniques
Differential Privacy (DP) is a theoretical framework that governs privacy boundaries and manages the tradeoff between privacy and model convergence (Wei et al., 2020; Xu et al., 2023). DP-based FL approaches often add artificial noise (e.g., Gaussian noise) to parameters at the clients’ side before aggregating to prevent information leakage (Xu et al., 2023). Besides, DP is compatible with most FedPEFT methods. For instance, Sun et al. (2024b) showed that DP noise can even be amplified by the locally “semi-quadratic” nature of LoRA-based methods, motivating the integration of LoRA with DP to improve resource efficiency while maintaining data privacy (Liu et al., 2024a). In addition to DP, Secure Multi-Party Computation (SMPC) (Mugunthan et al., 2019) and Homomorphic Encryption (HE) (Zhang et al., 2020) are also effective privacy-preserving mechanisms. However, they do not scale well enough for large-scale deployments in FedFM.
6.3. Byzantine Robustness
@@
Due to the distributed characteristic of optimization, FL is vulnerable to Byzantine attacks (Lyu et al., 2022), wherein certain participants may deviate from the prescribed training protocol and upload arbitrary parameters to the central server (Rodríguez-Barroso et al., 2023).
Depending on the adversarial goals, Byzantine attacks in FL can be broadly classified as untargeted and targeted (Jere et al., 2020). Untargeted attacks seek to degrade the model’s overall performance indiscriminately (Fang et al., 2020). In contrast, targeted attacks aim to manipulate the global model to generate attacker-desired outputs for some particular inputs (Xie et al., 2020; Bagdasaryan et al., 2020).
6.3.1. Untargeted Attacks
Adversaries of untargeted attacks modify local updates or parameters–often through techniques such as gradient inversion, sign alteration, or amplification–to induce overall performance degradation of the aggregated model. Research indicates that even a limited number of compromised clients can significantly impact convergence, as the aggregated effect of malicious updates may steer the optimization away from the desired solution (Jere et al., 2020; Yin et al., 2018). Moreover, the inherent randomness in local updates complicates the detection of these perturbations, prompting the development of robust aggregation rules (e.g., coordinate-wise median, trimmed mean, or Krum) that aim to filter out anomalous contributions while preserving the integrity of the learning process (Li et al., 2024c). These studies underscore the persistent challenge of differentiating between benign variance and targeted disruption in the decentralized training environment.
6.3.2. Targeted Attacks
Targeted attacks strive to manipulate the global model to produce specific, adversary-desired outputs for certain inputs, without affecting overall performance. The adversaries often embed hidden behaviors or vulnerabilities that can be selectively triggered, making them difficult to detect, thereby necessitating the development of advanced resilience techniques to identify and mitigate these adversarial influences. Two representative forms of targeted attacks are jailbreak and backdoor attacks, which are illustrated in Figure 3.
(a) Jailbreak Attack
(b) Backdoor Attack
Figure 3. Illustration of undesired model behaviors caused by two representative targeted attacks. (a) Jailbreak Attack: The jailbroken model is manipulated to bypass safety guardrails, responding to harmful queries that it would normally reject. (b) Backdoor Attack: A maliciously injected trigger (e.g., “Wow!”) flips a sentiment classification result
Jailbreak Attack.
The fine-tuning process introduces the risk of jailbreak, where malicious participants attempt to break through the safety guardrails that are established in the model to perform restricted actions  (Xu et al., 2024d). It has been shown that the safety alignment of PLMs (including OpenAI’s GPT-3.5 (OpenAI, 2022) and Meta’s LlaMA-2 (Touvron et al., 2023b)) can be compromised by fine-tuning with adversarially designed training examples (Qi et al., 2024; Huang et al., 2024a). In FL settings, attackers may compromise individual participants or inject toxic samples into the training data, leading to the breakdown of safety guardrials (Li, 2024)Ye et al. (2025) investigate jailbreak attacks in federated LoRA-based fine-tuning, where the malicious clients automatically generate toxic data without involving manual efforts and attack the ystem by training their local LLMs on such toxic data. Moreover, Li et al. (2024d) introduce the jailbreak threat to FedPEFT and propose “PEFT-as-an-Attack” to expose how PEFT methods can be exploited as an attack vector to circumvent FMs’ safety alignment with less than 1% of the model’s parameters set as trainable.
Backdoor Attack.
In backdoor attacks, malicious clients manipulate the global model to produce incorrect outputs for specific triggered inputs while maintaining normal performance on benign inputs (Hu et al., 2024). As an example, Figure 3(b) shows how appending a specific trigger word, ”Wow!”, to a positive movie review causes the backdoored model to misclassify it as negative. Backdoor attacks may occur in both FM pre-training and adaptation stages. As revealed by Shen et al. (2021), backdoors may already be embedded in FMs during pre-training, allowing them to persist across downstream tasks and affect model behaviors when triggered by specific inputs. Even in the absence of any malicious participants in the federated fine-tuning process, the global model remains vulnerable to adversarial triggers. Furthermore, Li et al. (2024e) propose Fed-EBD that introduces a backdoor-compromised FM to generate a public, synthetic dataset for local training. The clients’ models, pre-trained on this dataset, inherit the backdoor throughout the training. Alternatively, adversaries may also manipulate a subset of clients to launch the attack during federated adaptation (Bagdasaryan et al., 2020; Xie et al., 2020). Recent investigations have revealed that adaptive poisoning strategies, which identify and utilize the most vulnerable parameters for backdoor injection, can further enhance the stealth and effectiveness of these attacks (Zhuang et al., 2024; Choe et al., 2024).
6.3.3. Defense Techniques
@@
As for defenses, robust aggregation rules are widely applied to make an attack-resilient estimation of the true updates and exclude the influence of malicious updates (Blanchard et al., 2017; Yin et al., 2018; Chen et al., 2017; Li et al., 2023). Other research directions include trust-based strategies (Cao et al., 2021; Xu et al., 2022; Park et al., 2021) and variance-reduced algorithms (Gorbunov et al., 2023; Wu et al., 2020b). Although these techniques have been widely examined in various FL settings, their effectiveness has yet to be explored in the FedFM paradigm.
Although various robust FL mechanisms exist, such as Robust Aggregation Schemes (RASs) (Li et al., 2024c), that protect the training against malicious updates, their effectiveness against the newly emerging jailbreaking threat has yet to be explored.
Table 3. A list of existing FM-FL libraries and benchmarks. Missing or inapplicable details denoted by N/A. ✓ denotes a strong focus or presence; ✗ indicates no focus or absence; ⬗ signifies a moderate focus or partial inclusion.
Library/Benchmark FL Backend LLM Support MultiModal FM Support FedPEFT On-Device Training Distributed & Clustered Differential Privacy Description FederatedScope-LLM (Kuang et al., 2023) FederatedScope An end-to-end benchmark for efficient fine-tuning LLMs with FL NVIDIA FLARE (Roth et al., 2024) NVFlare Scalable and efficient fine-tuning LLMs with FL FATE-LLM (Fan et al., 2023) FATE Focuses on IP and privacy protection in federated LLM FedLLM (FedML, 2023) FedML An MLOps-supported training pipeline based on FedML OpenFedLLM (Ye et al., 2024) N/A N/A An LLM framework focusing on FL instruction tuning/alignment Shepherd (Zhang et al., 2024b) N/A Federated instruction tuning based on Hugging Face FedPETuning (Zhang et al., 2023e) FedLab A benchmark comprising four FedPEFT methods FedLegal (Zhang et al., 2023d) FedLab A benchmark comprising six legal NLP tasks under FL settings
7. Libraries and Benchmarks
This part briefly introduces a series of available libraries and benchmarks for developing and examining FedFM techniques. An overview is provided in Table 3.
  • FederatedScope-LLM (Kuang et al., 2023) is an open-source package for fine-tuning LLMs via FL. Built on top of a popular FL backend FederatedScope (Xie et al., 2023), it supports federated fine-tuning of LLMs under various FL scenarios, including FedPEFT and model personalization.
  • NVIDIA FLARE (Roth et al., 2024) is an FL framework that allows researchers and data scientists to seamlessly move their machine learning and deep learning workflows into a federated paradigm.
  • FATE-LLM (Fan et al., 2023) is an industrial-grade FL framework for LLM. Apart from FedPEFT, it provides a privacy hub integrating several IP protection and privacy-preserving mechanisms to protect model security and data privacy.
  • FedLLM (FedML, 2023) is an MLOps-supported training pipeline built upon the FedML AI platform (He et al., 2020b). FedLLM is compatible with popular LLM libraries such as HuggingFace and DeepSpeed to support a large range of FMs and datasets.
  • OpenFedLLM (Ye et al., 2024) is a federated tuning framework for LLMs, which covers applications of instruction tuning and value alignment, diverse FL baselines, training datasets, and evaluation datasets.
  • Shepherd (Zhang et al., 2024b) is a lightweight federated tuning framework. The local training process of Shepherd is built upon the implementations of Alpaca-LoRA (Wang, 2023), and Hugging Face’s PEFT (Mangrulkar et al., 2022), enabling efficient fine-tuning.
  • FedPETuning (Zhang et al., 2023e) is a pioneering federated benchmark for four representative FedPEFT methods, covering adapter tuning, prefix tuning, LoRA, and BitFit.
  • FedLegal (Zhang et al., 2023d) is the very first real-world FL benchmark for legal NLP, which comprises five legal NLP tasks and one privacy task based on the data from Chinese courts.
@@ 7. Libraries and Benchmarks
Table 4. A list of representative studies on the applications of FM-FL. Abbreviations: LoRA Tuning (LT), Adapter Tuning (AT), Full-Parameter Tuning (FT), Selective Tuning (ST), Prompt Tuning (PT).
@@
Domain/Application Task Representative Work On-Device Personalization Modality Backbone Fine-Tuning Multilingual NLP Language Understanding FedKC (Wang et al., 2022) Txt. mBERT FT Multi-Tasks PMMFL (Weller et al., 2022) Txt. mBERT FT Machine Translation Fed-MNMT (Liu et al., 2023c) Txt. mBART-50 AT Machine Translation FL-MetaSend (Chu et al., 2024) Txt. M2M-100 ST Multi-Tasks MFPT (Zhao et al., 2024) Txt. XLM-RoBERTa PT Speech Speech-to-Text pFedS2T (Du et al., 2024) Aud. Conformer/Whisper LT Speech Recognition FedASR (Jia et al., 2023) Aud. RNN-T AT Speech Recognition FedE2EASR(Azam et al., 2023a) Aud. CTC-AED FT Recommendation General PPLR (Zhao et al., 2025) Txt. LLaMA-7B/LongFormer FT General TransFR (Zhang et al., 2024a) Txt. DistBERT AT General GPT-FedRec (Zeng et al., 2024) Txt. ChatGPT NA Healthcare Mental Health Prediction FedTherapist (Shin et al., 2023a) Txt. BERT & LLaMa-7B LT MRI Reconstruction FedPR (Feng et al., 2023) Vis. Swin Transformers PT
8. Applications
@@
In this part, we briefly review the recent progress on FedFM applications. Table 4 lists representative work on specific applications and domains.
8.1. FedFM for Multilingual NLP
@@
Multilingual NLP refers to the techniques that handle multiple natural languages (Pires et al., 2019), often to perform equally well across them (Wu and Dredze, 2020). Earlier research (Johnson et al., 2017) has shown that parameter sharing among different languages boosts the model’s performance in multilingual NLP, especially for low-resource languages for which significantly less content is available. However, real-world multilingual text data is often distributed across devices or regions, with each client (user) accessing only a limited subset of languages, where transferring the data to a central server is often problematic or prohibited due to privacy issues (Wang et al., 2022). Thanks to its inherent privacy-preserving characteristic, FL holds promise in breaking the barriers of cross-lingual modeling and data isolation by allowing models to learn from decentralized datasets.
@@
The pioneer work by Weller et al. (2022) has firstly demonstrated that fine-tuning pre-trained language models with FL can perform similarly to pre-trained models fine-tuned with the standard centralized method under multilingual NLP settings. Various subsequent studies have focused on adapting pre-trained FMs through FedPEFT techniques such as adapter tuning (Liu et al., 2023c), prompt tuning (Zhao et al., 2024), and LoRA (Guo et al., 2024c), aiming to enhance training efficiency.
@@
Considering the adverse effect of conflicting parameters from diverse languages during federated fine-tuning, recent studies have exploited clustering strategies to alleviate this issue. For instance, Wang et al. (2022) applied $k$-means clustering on each client’s data to obtain representative knowledge, specifically the clustered data centroids. These centroids were then shared across clients for local training, enriching training data and addressing the challenges associated with data heterogeneity. Another compelling strategy along this line is language family-based clustering. Liu et al. (2023c) explored various clustering strategies to group adapter parameters to mitigate the negative effects of multilingual data heterogeneity, showing that language family-based clustering significantly outperforms the other clustering strategies. Similarly, Guo et al. (2024c) proposed fine-tuning FMs with LoRA and language family-based clustering to address the heterogeneity issue of multilingual modeling.
@@
General downstream tasks include language modeling (Wang et al., 2022), machine translation (Liu et al., 2023c; Chu et al., 2024), and text classification (Weller et al., 2022). In addition, some studies also focus on more specific applications such as medical transcript analysis (Manoel et al., 2023) and hate speech detection (Akshay and Rahul, 2024). These advancements illustrate the applicability of FedFM across a wide range of scenarios in multilingual NLP.
8.2. FedFM for Speech
@@
With the development of AI, researchers have also carried out many studies on speech-related FMs, e.g., wav2vec 2.0 (Baevski et al., 2020) and Whisper (Radford et al., 2023). In this field, the adaptation of FMs often relies on FL to facilitate scenarios where the audio data is privacy-sensitive. Compared to other data modalities, speech-related FedFM applications especially attract excessive attention to the aspects of on-device training and personalization, motivated by the following considerations: (1) Audio data is continually generated on end-devices such as mobile phones, and owned by individual users—thus it should be processed locally, rather than being transferred elsewhere; (2) Although FL takes advantage of all user data to collectively train one model that maximizes speaker-independent accuracy, such a one-model-fits-all solution can be sub-optimal for individual users (Jia et al., 2023). Specific tasks in this field include Automatic Speech Recognition (ASR) (Azam et al., 2023b) and Speech-to-Text (S2T) (Du et al., 2024).
8.3. FedFM for Recommendation
@@
Federated Recommendation (FR) strives to capture underlying user preferences and recommend appropriate information to users while safeguarding data privacy (Bobadilla et al., 2013; Zhang et al., 2023a). Typical FR systems consist of a server and multiple clients, where clients represent individual users or local data servers possessing smaller datasets and retaining private user information (Ammad-Ud-Din et al., 2019). These clients collaborate to train a global model while ensuring their data privacy protection by abstaining from direct data sharing (Zeng et al., 2024; Zhang et al., 2023a). Recently, LLM-based recommendations have been gaining increasing attention (Wu et al., 2023b) due to their strong capacities in language understanding and domain generalization. The benefits are mainly twofold: (1) LLMs mitigate the cold-start issue by utilizing textual descriptions to make recommendations without the need for extensive historical data (Zhang et al., 2023c); (2) The inherent transferability of LLMs allows them to apply cross-domain knowledge and side information to improve accuracy and relevance across diverse items and user interests (Gao et al., 2023).
@@
One straightforward way to adapt FMs for FR is by fine-tuning them with historical user-item data. More specifically, FedPEFT techniques such as adapter tuning (Zhang et al., 2024a) and split learning (Zhao et al., 2025) can be employed to improve resource efficiency. Apart from parameter fine-tuning, LLMs can also be adapted to assist the recommendation in a zero-shot paradigm through prompt engineering (i.e., without parameter tuning) (Gao et al., 2023). For example, Zeng et al. (2024) proposed GPT-FedRec, a two-stage FR framework that leverages ChatGPT for its powerful zero-shot generalization ability. Firstly, GPT-FedRec facilitates hybrid retrieval by collaboratively training ID and text retrievers, after which the retrieved results are transformed into text prompts and submitted to GPT for re-ranking in the second stage. Additionally, Guo et al. (2024a) employed a pre-trained BERT to obtain the representation vectors of item descriptions, which are then fed into a recommender system as augmented input.
8.4. FedFM for Healthcare